What is the maximum power that can be delivered by a solid propeller shaft of diameter 50 millimeters if the shaft is turning at 600 revolutions per minute and the shaft material has a shear modulus of elasticity of 80 GPa and an allowable shear stress is 100 MPa?
Answer #1

Answer)

Diameter of shaft = 50mm

Let shear stress = S = 100 MPa = 100 N/m2

Torque acting on shaft = T

Polar moment of inertia = J = (πd^4)/32

r = radius of shaft

From the equation

S/r = T/J  

We get

(100/25) = T/(πd^4)/32

T = 4*(π*50^4)/32

T = 24,53,125 Nmm

T = 2.45 KNM

Power delivered by shaft = 2πNT/60

Here N = 600 revolutions/min

P = (2π*600*2.45)/60 = 153.8 KW

Power delivered is 153.8 kilowatts